OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (anonymous):

ok a bit long but, you have to solve both equations more then once, 4x+y=11 becomes y = 11 -4x x+2y=8 becomes x = 8-2y now you substitute the values of x , y in the original equations, 4(x) +y = 11 4(8-2y) +y = 11 distribute the 4 32 - 8y +y = 11 subtract 32 from both sides for -8y + y = -21 add -8y and y together for -7y -7y = -21 multiply by -1 to make positive and divide both sides by 7 for y = 3 then do the same for x